OpenStudy (anonymous):

Write the standard equation of the circle with center (-3,5) && diameter of 10. A. (X+3) + (Y-5)^2 = 25 B. (X+3) + (Y-5)^2 = 100 C. (X-5) + (Y+3)^2 = 5 D. (X-3) + (Y+5)^2 = 85

OpenStudy (anonymous):

(x+3)^2+(y-5)^2=25

OpenStudy (anonymous):

how did you get that ? @Federica

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@apoorvk

OpenStudy (apoorvk):

Hmm.. for a circle with center at (a,b) and radius 'r' the equation is: \[(x-a)^2 + (y-b)^2 = r^2\] Now you have the coordinates of the center, as well as the diameter, from which you can find the radius. Helps?

OpenStudy (anonymous):

Yes, you´re right apoorvk.

OpenStudy (anonymous):

ok well i plugged in the solution for A but -3 + 3 = 0 && 5+ -5 = 0

OpenStudy (anonymous):

@apoorvk . i think its B

OpenStudy (apoorvk):

you don't even need to plug in, just substitute values for 'a', 'b' and 'r' in the equation formula.

OpenStudy (apoorvk):

umm, slight bit of a mistake @Senior012 , the *diameter* is 10, but you need to substitute the *radius* in there.

OpenStudy (anonymous):

sorry i was looking at a different question.. C ?

OpenStudy (anonymous):

wait D ?

OpenStudy (apoorvk):

well, radius = 10/2 = 5 so, r^2 = 5^2 = 25 put this in place of r^2

OpenStudy (apoorvk):

Lol, stay calm, you're getting confused! Cool man :)

OpenStudy (anonymous):

ok so my guess is A && C ? right or wrong ? nothing else makes sense

OpenStudy (apoorvk):

C makes no sense, since r=5, but r^2 = 25. (Have a look at the formula again.) so that leaves us with A!

OpenStudy (apoorvk):

helps eh?

OpenStudy (anonymous):

But why the term of x is not squared?

OpenStudy (apoorvk):

Hmm.. good point, I didn't even notice it.. Is that a typo @Senior012 ?